OpenStudy (texaschic101):

I know this is a stupid question....but I am having a problem :

OpenStudy (texaschic101):

In a box and whisker plot, can Q3 be smaller then the median (Q2) ?

OpenStudy (anonymous):

no it cant :)

OpenStudy (texaschic101):

Look at this.... data : 10,10,15,19,25 Q1 = 20/2 = 10 Q2 = 15 Q3 = 44/2 = 11 lower extreme = 10 upper extreme = 25 what am I doing wrong ?

OpenStudy (texaschic101):

oops.....I feel stupid.....nvm....I see what I did.....Q3 = 22.